The freezing point of ethanol is –114.6 oC. The molal freezing point depression constant for ethanol is 2.00 oC/m. What is the freezing point (oC) of a solution prepared by dissolving 170.0 g of glycerin (C3H8O3, a nonelectrolyte) in 250 g of ethanol?
1
Expert's answer
2020-05-13T14:20:04-0400

"\u2206T_f=K_fm"

Given : "K_f= 2\u00b0C\/m; T_f=-114.6\u00b0C" (actual freezing point)

m=moles of solute /mass of solvent

"m=170\/(0.25*92)=7.3913molal"

"\u2206T_f=7.3913*2=14.7826\u00b0C"

Thus apparent freezing point is

"T_f'=T_f-14.7826=-128.8426\u00b0C"